# Higgs Phases and Boundary Criticality in memoriam Peter Higgs

### v1

@@ III.6 Summary and Discussion: Non-Abelian Case
In this section we have discussed global boundary symmetries in non-Abelian Higgs models and their spontaneous symmetry breaking with both fundamental Higgs fields and group-valued Higgs fields. For general non-Abelian gauge group $\mathcal{G}$ there is a global $\mathcal{G}$ symmetry acting on the boundary of the system, which is equivalent by the Gauss law to the total color charge of the system, Eq. 48. For the group-valued case, there is an additional bulk matter symmetry given by right multiplication of the matter field. If $\mathcal{G}$ is an Abelian group, these two symmetries exactly coincide, since there is no distinction between left and right multiplication. For $\mathcal{G}=\mathrm{SU}(2)$, the two types of models are equivalent due to fact that $\mathrm{SU}(2)$ is equivalent to the unit quaternions. For other gauge groups the two types of models are inequivalent. In all cases, the location of the boundary transition that we identify can be shifted by tuning the boundary coupling.
@@
In the group-valued case, the infinite $\kappa$ limit freezes the bulk of the system. The resulting boundary theory can be expressed in terms of gauge-invariant half-open Wilson strings at the boundary, Eq. 60. This may be viewed as a lattice discretization of a principal chiral model, with $\mathcal{G}\times\mathcal{G}$ symmetry coming from the boundary color flux and bulk matter symmetries. This global symmetry is expected to be broken to the diagonal subgroup at large $\beta$. Since the boundary is frozen in the $\cramped{\kappa\to\infty}$ limit and should remains strongly gapped at large $\kappa$, the boundary transition is expected to persist also for finite $\kappa$. Our SU(2) Monte Carlo simulations verify this, and the boundary transition line appears to terminate at the bulk critical endpoint.
@@
For Higgs fields in the fundamental representation, there is only a single global $\mathcal{G}$ symmetry, generated by the total color charge of the system and acting on the boundary. The general picture is that the bulk $\mathcal{G}$ gauge symmetry is Higgsed down to a subgroup $\mathcal{H}$. If $\mathcal{H}$ is trivial, then the bulk is completely frozen in the $\cramped{\kappa\to\infty}$ limit. If $\mathcal{H}$ is Abelian, there may be a bulk phase transition even at infinite $\kappa$ (though that does not preclude a boundary symmetry breaking). If $\mathcal{H}$ is non-Abelian then the bulk reduces to a gapped $\mathcal{H}$ gauge theory as $\cramped{\kappa\to\infty}$. The boundary theory is expected to be a gauged nonlinear $\sigma$-model with target space $\mathcal{G}/\mathcal{H}$, Eq. 74. Note that by placing the Higgs in other representations (e.g. adjoint) or adding addition Higgs fields, one may target different subgroups $\mathcal{H}$  [44] and thus obtain different boundary theories. For example, by starting from an $\mathrm{SO}(N)$ gauge theory with $M$ Higgs fields in the fundamental representation, the general expectation is that it reduces to an $\mathrm{SO}(N-M)$ gauge theory in the bulk, and the boundary target space is the Stiefel manifold $\mathrm{SO}(N)/\mathrm{SO}(N-M)$, which have recently attracted significant interest [45]
@@
We considered in particular the case $\mathrm{SU}(N)$ (and by direct extension SO($N$)). In the $\cramped{\kappa\to\infty}$ limit, the bulk Higgses down to a $\mathrm{SU}(N-1)$ gauge theory, while the boundary decomposes into a pure $\mathrm{SU}(N-1)$ part and an $\mathrm{SU}(N)$ rotor part. The global symmetry, under most circumstances, will be broken spontaneously at large $\beta$. We argued that this SU($N$) fundamental Higgs boundary phase transition lies in the O($2N$) universality class, since there are no relevant operators differentiating $\mathrm{SU}(N)$ from $O(2N)$. We tested this prediction numerically for the $\mathrm{SU}(3)$ lattice gauge theory with fundamental Higgs, demonstrating a clean scaling collapse with O(6)6)6 ) critical exponents.
@@
One interesting exceptional case is SO($3$) gauge theory with fundamental Higgs. This is equivalent to the Georgi-Glashow electroweak theory, with gauge group $\mathrm{SU}(2)$ and the Higgs field in the adjoint representation. In the $\kappa\to\infty$ limit we can fix to unitary gauge which Higgses the bulk gauge group down to $\mathrm{SO}(2)\simeq\mathrm{U}(1)$. In $D=4$ and $\kappa\rightarrow\infty$, the bulk has a confinement-deconfinement transition as a function of $\beta$, while we predict an additional O($3$) boundary critical point in the Higgs regime. This may have implications for the physics of domain walls or cosmic strings formed in the early universe.
@@
IV Boundary Symmetry Breaking in Higher-Form Abelian Higgs Models
@@
IV.1 Higher-Form Abelian Higgs Models
[…]
@@ IV.1 Higher-Form Abelian Higgs Models
where the sum is over the $k$-cells forming the oriented boundary of the $(k+1)$-cell $c$. This is a straightforward generalization of the Abelian Higgs model, Eq. 8, to a theory of extended $(k-1)$-dimensional charged objects attached to $k$-dimensional electric flux branes [46].
@@
for an arbitrary $(k-1)$-form $\lambda$. We will see that this enforces the Gauss law attaching electric branes to the charged objects. When $k>1$, this gauge invariance includes “gauge-of-gauge” transformations
@@
for arbitrary $(k-2)$-form $\alpha$. This enforces an additional Gauss law which states that the $(k-1)$-dimensional electrically charged objects are closed.
When $\kappa=0$, Eq. 79 reduces to a pure $k$-form U(1) gauge theory, which has a global electric $k$-form symmetry given by
@@
In the limit $\beta\to\infty$ it has a $k_{\mathrm{m}}$-form magnetic symmetry, discussed further in Section IV.2.
The phase diagram of this model is expected to be similar to that of the 1-form U(1) gauge theory, sketched in Fig. 1, as long as $\cramped{D>k+2}$ ($\cramped{k_{\text{m}}>0}$). In the marginal case $\cramped{D=k+2}$ ($\cramped{k_{\text{m}}=0}$), the magnetic defect is an instanton (zero dimensional in spacetime) and is expected to destabilize the deconfined phase [55, 56, 57]—a generalization of the Polyakov mechanism for 1-form U(1) gauge theory in $\cramped{D=3}$ [58], which itself may be viewed as a higher-form generalization of the Mermin-Wagner theorem for 0-form symmetries [3, 4]. In those marginal cases, the bulk phase diagram will generically be similar to that of the non-Abelian models, as in Fig. 5 [55, 59, 60].
IV.1.1 Gauss Law and Higher-Form Matter Symmetry
@@
The Hamiltonian formulation follows the exact arguments laid out in Section II.1, with conjugate operators $[\hat{\theta},\hat{n}]=i$ on all $(k-1)$-cells and $[\hat{A},\hat{E}]=i$ on all $k$-cells. Invariance of physical states under the gauge transformations, Eq. 81, enforces the Gauss law
$-(\mathrm{d}^{\dagger}\hat{E})_{c}\equiv-\sum_{\mathclap{c^{\prime}\in\partial^{\dagger}c}}\hat{E}_{c^{\prime}}=\hat{n}_{c}\quad(c\in X_{k-1}),$ (85)
where $c$ is a $\cramped{(k-1)}$-cell and the sum is over its coboundary, the set of $k$-cells $c^{\prime}$ containing $c$ in their positively-oriented boundary. This simply says that the number of electric $k$-branes emanating from $c$ is equal to the amount of electric charge on $c$. For $\cramped{k>1}$, the charges carry a sense of orientation. For example, when $\cramped{k=2}$ the charges are nothing but the electric strings of a 1-form gauge field. The “gauge-of-gauge” symmetry, Eq. 82, enforces the constraint
@@
which says that the charged objects are closed.666 This constraint obviously follows from the Gauss law Eq. 85 (following from $\cramped{\mathrm{d}^{2}=0}$), just as Eq. 82 is already implied by Eq. 81, but it is worth spelling out for those unfamiliar with this point.
In the $k=1$ theory discussed in Section II with 0-form Higgs field, the net charge of the Higgs field generates a global 0-form symmetry, which is pure gauge with periodic boundaries but becomes physical with the choice of electric boundary conditions. The natural extension of this global matter symmetry for general $k$ is a $\cramped{(k-1)}$-form symmetry,
@@
where $c$ is the $\cramped{(k-1)}$-cell in the direct lattice corresponding to $\tilde{c}$ in the dual lattice, and $\delta_{\tilde{\Sigma}}$ is a $\cramped{(k-1)}$-form Poincaré dual to $\tilde{\Sigma}$.777 For our purposes, the defining property of the Poincaré dual of a $\cramped{(d-k)}$-dimensional closed surface $\tilde{\Sigma}$ in the dual lattice is that it is a $k$-form in the direct lattice acting as a generalized delta function, for example the unit $k$-form supported on the direct lattice $k$-cells piercing $\tilde{\Sigma}$ with the appropriate orientation. Note that $\delta_{\tilde{\Sigma}}$ is only defined up to an exact form because $\tilde{\Sigma}$ is closed, i.e. it is a cohomology class. These operators generate the symmetry, Eq. 87,
@@
These operators are topological, i.e. they only depend on the homology class of $\tilde{\Sigma}$, owing to the matter Gauss law Eq. 86. Consider replacing it with another surface such that $\cramped{{\tilde{\Sigma}}^{\prime}-\tilde{\Sigma}=\partial\tilde{V}}$ for some $\cramped{(d-(k-2))}$-volume $\tilde{V}$. Then $\cramped{\delta_{\tilde{\Sigma}^{\prime}}=\delta_{\tilde{\Sigma}}+\mathrm{d}\delta_{\tilde{V}}}$. Plugged into Eq. 88, we have
$\hat{Q}(\tilde{\Sigma}^{\prime})=\langle\hat{n},\delta_{\tilde{\Sigma}}\rangle+\langle\mathrm{d}^{\dagger}\hat{n},\delta_{\tilde{V}}\rangle=\hat{Q}(\tilde{\Sigma}),$ (90)
where we used the matter Gauss law, Eq. 86. Therefore there is one $\cramped{(k-1)}$-form charge generator for each homology class in $H_{d-(k-1)}$. Deforming $\tilde{\Sigma}$ without changing its homology class corresponds to shifting $\theta$ by an exact form, which are just the gauge transformations of Eq. 82.
The operators charged under $\hat{Q}(\tilde{\Sigma})$ are the charge creation/annihilation operators, i.e. Wilson branes supported on open $k$-dimensional surfaces $M$ which insert an electric membrane with charge on its boundary,
$e^{i\alpha\hat{Q}(\tilde{\Sigma})}e^{i(\mathrm{d}\hat{\theta}-A)(M)}$ $e^{-i\alpha\hat{Q}(\tilde{\Sigma})}$
$=e^{-i\hat{A}(M)}e^{i\alpha\hat{Q}(\tilde{\Sigma})}e^{i\hat{\theta}(\partial M)}e^{-i\alpha\hat{Q}(\tilde{\Sigma})}$
@@
#$\#(\partial M,\tilde{\Sigma})=\delta_{\tilde{\Sigma}}(\partial M).$ (92)
Thus the operator $\hat{Q}(\tilde{\Sigma})$ simply counts the intersection number of the closed charged objects with $\tilde{\Sigma}$.
In the Maxwell case, $\cramped{k=1}$, the matter charges are point particles, $\tilde{\Sigma}$ is a closed $d$-dimensional volume in the dual lattice, and the only non-trivial choice is to take it to be all of space. The associated charge then simply counts the number of positive minus the number of negative charges. For a less trivial example, consider the case $\cramped{k=2}$ and $\cramped{d=3}$, so that the matter charges are 1-dimensional strings and $\tilde{\Sigma}$ is a closed two-dimensional surface in the dual lattice, intersecting a collection of links in the direct lattice. A choice of its Poincaré dual is a 1-form which is zero everywhere except on the intersected links, on which it has unit value in the direction normal to the surface. The charge operator, Eq. 88, then counts (with signs, because $\cramped{\hat{n}_{-c}=-\hat{n}_{c}}$) the number of strings piercing the surface $\tilde{\Sigma}$, i.e. the intersection number between closed strings and the surface. Because the charges are closed strings, this intersection number must be zero unless $\tilde{\Sigma}$ winds around a periodic boundary and intersects a charge which also winds around the periodic boundaries in the transverse direction.
In a closed system, gauge invariance guarantees that the matter charge operators are exactly zero because of the Gauss law, Eq. 85, which when inserted into Eq. 88 yields $\cramped{\langle\hat{E},\mathrm{d}\delta_{\tilde{\Sigma}}\rangle=0}$. Equivalently, the intersection numbers, Eq. 92, are exactly zero because $\partial M$ has trivial homology class, i.e. $\cramped{\delta_{\tilde{\Sigma}}(\partial M)=\mathrm{d}\delta_{\tilde{\Sigma}}(M)=0}$ since $\tilde{\Sigma}$ is closed. Consider for example the case $k=2$, where the charge operators count the number of charged strings wrapped around a periodic boundary. We cannot, in a closed system, have a single non-contractible charged string, because it is attached to an electric membrane which must end somewhere inside the system. The only possibility is that it ends on another non-contractible charged string going the other direction, such that the two strings constitute the boundary of the electric membrane. This is the sense in which a closed system is charge-neutral when $k>1$.
@@
Figure 10: (a) An illustration of the relation between the direct lattice and dual lattice at the boundary of the system in $d=3$ spatial dimensions. The direct lattice is colored the same as in Fig. 2. The dual lattice is indicated by black dual sites, white dual edges, and purple dual plaquettes. Each $k$-cell of the direct lattice corresponds to a $\cramped{(d-k)}$-cell in the dual lattice. Because the boundary ($\partial X$) of the direct lattice ($X$) is open with cells “sticking out” (green), the dual boundary ($\partial\tilde{X}$) is closed and has no protruding cells. We also color the $(D-1)$-dimensional edge layer ($Y$) dark gray, which is where the $\kappa=\infty$ boundary action is defined, Eq. 100. (b) An illustration of a termination of $\tilde{\Sigma}$ on the boundary for a 1-form Higgs field coupled to a 2-form gauge field in $d=3$ spatial dimensions. $\tilde{\Sigma}$ is a two-dimensional surface (cyan) which terminates at the boundary along a 1-dimensional contour (green). The Gauss law makes $\hat{Q}(\tilde{\Sigma})$ equal to the sum of the electric fluxes $\hat{E}_{p}$ on the boundary plaquettes pierced by the edge of $\tilde{\Sigma}$ (green squares with solid edges), Eq. 94. The edge $\partial\tilde{\Sigma}$ then intersects the half-open Wilson membranes attached to a charged string (red curve) attached to an electric membrane (yellow) exiting the system through the open boundary.
IV.1.2 Open Boundaries and Global Matter Symmetry
To make the charge operators non-trivial, we must impose boundary conditions such that either $\tilde{\Sigma}$ can terminate on the boundary, and $\delta_{\tilde{\Sigma}}$ is not a closed form, or such that the contour $\partial M$ can terminate at the surface and thus not be closed. The latter case, however, means that the charged objects can pass through the boundary, meaning charge is not conserved. Therefore, extending our results from Section II, we take the former case. We consider open boundaries in the same form as Fig. 2. To be precise, the boundary consists of a layer of $D$-dimensional hypercubes, with all cells in the boundary layer which do not touch a bulk cell removed (or, equivalently, $A=0$ and $\theta=0$ on those cells). In other words, we remove the cells on the vacuum side, creating links missing one end, plaquettes missing one edge, cubic cells missing one face, etc. The boundary action is given by

### v2

@@ 3.6 Summary and Discussion: Non-Abelian Case
In this section we have discussed global boundary symmetries in non-Abelian Higgs models and their spontaneous symmetry breaking with both fundamental Higgs fields and group-valued Higgs fields. For general non-Abelian gauge group $\mathcal{G}$ there is a global $\mathcal{G}$ symmetry acting on the boundary of the system, which is equivalent by the Gauss law to the total color charge of the system, Eq.˜48. For the group-valued case, there is an additional bulk matter symmetry given by right multiplication of the matter field. If $\mathcal{G}$ is an Abelian group, these two symmetries exactly coincide, since there is no distinction between left and right multiplication. For $\mathcal{G}=\mathrm{SU}(2)$, the two types of models are equivalent due to fact that $\mathrm{SU}(2)$ is equivalent to the unit quaternions. For other gauge groups the two types of models are inequivalent. In all cases, the location of the boundary transition that we identify can be shifted by tuning the boundary coupling.
@@
In the group-valued case, the infinite $\kappa$ limit freezes the bulk of the system. The resulting boundary theory can be expressed in terms of gauge-invariant half-open Wilson strings at the boundary, Eq.˜60. This may be viewed as a lattice discretization of a principal chiral model, with $\mathcal{G}\times\mathcal{G}$ symmetry coming from the boundary color flux and bulk matter symmetries. This global symmetry is expected to be broken to the diagonal subgroup at large $\beta$. Since the boundary is frozen in the $\cramped{\kappa\to\infty}$ limit and should remains strongly gapped at large $\kappa$, the boundary transition is expected to persist also for finite $\kappa$. Our SU(2) Monte Carlo simulations verify this, and the boundary transition line appears to terminate at the bulk critical endpoint.
@@
For Higgs fields in the fundamental representation, there is only a single global $\mathcal{G}$ symmetry, generated by the total color charge of the system and acting on the boundary. The general picture is that the bulk $\mathcal{G}$ gauge symmetry is Higgsed down to a subgroup $\mathcal{H}$. If $\mathcal{H}$ is trivial, then the bulk is completely frozen in the $\cramped{\kappa\to\infty}$ limit. If $\mathcal{H}$ is Abelian, there may be a bulk phase transition even at infinite $\kappa$ (though that does not preclude a boundary symmetry breaking). If $\mathcal{H}$ is non-Abelian then the bulk reduces to a gapped $\mathcal{H}$ gauge theory as $\cramped{\kappa\to\infty}$. The boundary theory is expected to be a gauged nonlinear $\sigma$-model with target space $\mathcal{G}/\mathcal{H}$, Eq.˜75. Note that by placing the Higgs in other representations (e.g. adjoint) or adding addition Higgs fields, one may target different subgroups $\mathcal{H}$  [52] and thus obtain different boundary theories. For example, by starting from an $\mathrm{SO}(N)$ gauge theory with $M$ Higgs fields in the fundamental representation, the general expectation is that it reduces to an $\mathrm{SO}(N-M)$ gauge theory in the bulk, and the boundary target space is the Stiefel manifold $\cramped{\mathrm{SO}(N)/\mathrm{SO}(N-M)}$, which have recently attracted significant interest [53]
@@
We considered in particular the case $\mathrm{SU}(N)$ (and by direct extension $\mathrm{SO}(N)$). In the $\cramped{\kappa\to\infty}$ limit, the bulk Higgses down to a $\mathrm{SU}(N-1)$ gauge theory, while the boundary decomposes into a pure $\mathrm{SU}(N-1)$ part and an $\mathrm{SU}(N)$ rotor part. The global symmetry, under most circumstances, will be broken spontaneously at large $\beta$. We argued that this $\mathrm{SU}(N)$ fundamental Higgs boundary phase transition lies in the $\mathrm{O}(2N)$ universality class, since there are no relevant operators differentiating $\mathrm{SU}(N)$ from $\mathrm{O}(2N)$. We tested this prediction numerically for the $\mathrm{SU}(3)$ lattice gauge theory with fundamental Higgs, demonstrating a clean scaling collapse with $\mathrm{O}(6)$ critical exponents.
@@
One interesting exceptional case is $\mathrm{SO}(3)$ gauge theory with fundamental Higgs. This is equivalent to the Georgi-Glashow electroweak theory, with gauge group $\mathrm{SU}(2)$ and the Higgs field in the adjoint representation. In the $\kappa\to\infty$ limit we can fix to unitary gauge which Higgses the bulk gauge group down to $\mathrm{SO}(2)\simeq\mathrm{U}(1)$. In $D=4$ and $\kappa\rightarrow\infty$, the bulk has a confinement-deconfinement transition as a function of $\beta$, while we predict an additional $\mathrm{O}(3)$ boundary critical point in the Higgs regime. This may have implications for the physics of domain walls or cosmic strings formed in the early universe.
@@
4 Boundary Symmetry Breaking in Higher-Form Abelian Higgs Models
@@
4.1 Higher-Form Abelian Higgs Models
Here we consider the case of $k$-form U(1) gauge theory, though restriction to $\mathbbm{Z}_{N}$ subgroups follows. Let $A$ be a $k$-form field and $\theta$ a (k1)(k-1)-form field, each taking values in $\mathbbm{R}/2\pi\mathbbm{Z}$ in $D$ Euclidean spacetime dimensions, with $\cramped{1\leq k\leq D-2}$.555 If $k=0$ there is no gauge field. If k=D1k=D-1 then the gauge field is not dynamical and can be completely integrated out using the Gauss law, yielding long-range interactions for the Higgs field. If $k=D$ then $\cramped{\mathrm{d}A=0}$ identically. By a $k$-form we mean a function $\omega$ on oriented $k$-dimensional cells $c$ of the lattice, such that $\omega(-c)=-\omega(c)$, where c-c denotes the cell $c$ with the opposite orientation. See Appendix˜A for a more detailed discussion of the discrete differential forms notation used throughout this section. Let $X_{k}$ denote the collection of $k$-cells of the lattice, each with a fixed orientation. The fields are governed by the generalized Fradkin-Shenker action
@@ 4.1 Higher-Form Abelian Higgs Models
where the sum is over the $k$-cells forming the oriented boundary of the $(k+1)$-cell $c$. This is a straightforward generalization of the Abelian Higgs model, Eq.˜8, to a theory of extended (k1)(k-1)-dimensional charged objects attached to $k$-dimensional electric flux branes [54].
@@
for an arbitrary (k1)(k-1)-form $\lambda$. We will see that this enforces the Gauss law attaching electric branes to the charged objects. When $k>1$, this gauge invariance includes “gauge-of-gauge” transformations
@@
for arbitrary (k2)(k-2)-form $\alpha$. This yields an additional Gauss-type constraint which enforces that the (k1)(k-1)-dimensional electrically charged objects are closed. [^6] In the continuum formulation these extra constraints are the pure-spatial components of the conserved higher-form Noether current [23].
When $\kappa=0$, Eq.˜81 reduces to a pure $k$-form U(1) gauge theory, which has a global electric $k$-form symmetry given by
@@
In the limit $\cramped{\beta\to\infty}$ it has a $k_{\mathrm{m}}$-form magnetic symmetry, discussed further in Section˜4.2.
The phase diagram of this model is expected to be similar to that of the 1-form U(1) gauge theory, sketched in Fig.˜1, as long as $\cramped{D>k+2}$ ($\cramped{k_{\text{m}}>0}$). In the marginal case $\cramped{D=k+2}$ ($\cramped{k_{\text{m}}=0}$), the magnetic defect is an instanton (zero dimensional in spacetime) and is expected to destabilize the deconfined phase [63, 64, 65]—a generalization of the Polyakov mechanism for 1-form U(1) gauge theory in $\cramped{D=3}$ [66], which itself may be viewed as a higher-form generalization of the Mermin-Wagner theorem for 0-form symmetries [4, 5]. In those marginal cases, the bulk phase diagram should be qualitatively similar to that of the non-Abelian models as in Fig.˜5 [63, 67, 68].
4.1.1 Gauss Law and Matter Symmetry
@@
The Hamiltonian formulation follows the exact arguments laid out in Section˜2.1, with conjugate operators $\hat{\theta},\hat{n}]=i$ on all (k1)(k-1)-cells and $\hat{A},\hat{E}]=i$ on all $k$-cells. Invariance of physical states under the gauge transformations, Eq.˜83, enforces the Gauss law
$(\mathrm{d}^{\dagger}\hat{E})_{c}\equiv-\sum_{\mathclap{c^{\prime}\in\partial^{\dagger}c}}\hat{E}_{c^{\prime}}=\hat{n}_{c}\quad(c\in X_{k-1}),$ (87)
where $c$ is a $\cramped{(k-1)}$-cell and the sum is over its coboundary, the set of $k$-cells $c^{\prime}$ containing $c$ in their positively-oriented boundary. This simply says that the number of electric $k$-branes emanating from $c$ is equal to the amount of electric charge on $c$. For $\cramped{k>1}$, the charges carry a sense of orientation. For example, when $\cramped{k=2}$ the charges are nothing but the electric strings of a 1-form gauge field. The “gauge-of-gauge” symmetry, Eq.˜84, enforces the constraint
@@
which says that the charged objects are closed.777 This constraint obviously follows from the Gauss law Eq. 87 (following from $\cramped{\mathrm{d}^{2}=0}$), just as Eq. 84 is already implied by Eq. 83, but it is worth spelling out for those unfamiliar with this point.
In the $k=1$ theory discussed in Section˜2 with 0-form Higgs field, the net charge of the Higgs field generates a global 0-form symmetry, which is pure gauge with periodic boundaries but becomes physical with the choice of electric boundary conditions. The natural extension of this global matter symmetry for general $k$ is a $\cramped{(k-1)}$-form symmetry,
@@
where $c$ is the $\cramped{(k-1)}$-cell in the direct lattice corresponding to $\tilde{c}$ in the dual lattice, and $\delta_{\tilde{\Sigma}}$ is a $\cramped{(k-1)}$-form Poincaré dual to $\tilde{\Sigma}$.888 For our purposes, the defining property of the Poincaré dual of a $\cramped{(d-k)}$-dimensional closed surface $\tilde{\Sigma}$ in the dual lattice is that it is a $k$-form in the direct lattice acting as a generalized delta function, for example the unit $k$-form supported on the direct lattice $k$-cells piercing $\tilde{\Sigma}$ with the appropriate orientation. Note that $\delta_{\tilde{\Sigma}}$ is only defined up to an exact form because $\tilde{\Sigma}$ is closed, i.e. it is a cohomology class. These operators generate the symmetry, Eq.˜89,
@@
These operators are topological, i.e. they only depend on the homology class of $\tilde{\Sigma}$, owing to the matter Gauss law Eq.˜88. Consider replacing it with another surface such that $\cramped{{\tilde{\Sigma}}^{\prime}-\tilde{\Sigma}=\partial\tilde{V}}$ for some $\cramped{(d-(k-2))}$-volume $\tilde{V}$. Then $\cramped{\delta_{\tilde{\Sigma}^{\prime}}=\delta_{\tilde{\Sigma}}+\mathrm{d}\delta_{\tilde{V}}}$. Plugged into Eq.˜90, we have
[…]
where we used the matter Gauss law, Eq.˜88. Therefore there is one $\cramped{(k-1)}$-form charge generator for each homology class in $H_{d-(k-1)}$. Deforming $\tilde{\Sigma}$ without changing its homology class corresponds to shifting $\theta$ by an exact form, which are just the gauge transformations of Eq.˜84.
[…]
$e^{i\alpha\hat{Q}(\tilde{\Sigma})}e^{i(\mathrm{d}\hat{\theta}-A)(M)}e^{-i\alpha\hat{Q}(\tilde{\Sigma})}$ $=e^{-i\hat{A}(M)}e^{i\alpha\hat{Q}(\tilde{\Sigma})}e^{i\hat{\theta}(\partial M)}e^{-i\alpha\hat{Q}(\tilde{\Sigma})}$
@@
$(\partial M,\tilde{\Sigma})=\delta_{\tilde{\Sigma}}(\partial M).$ (94)
@@
In the Maxwell case, $\cramped{k=1}$, the matter charges are point particles, $\tilde{\Sigma}$ is a closed $d$-dimensional volume in the dual lattice, and the only non-trivial choice is to take it to be all of space. The associated charge then simply counts the number of positive minus the number of negative charges. For a less trivial example, consider the case $\cramped{k=2}$ and $\cramped{d=3}$, so that the matter charges are 1-dimensional strings and $\tilde{\Sigma}$ is a closed two-dimensional surface in the dual lattice, intersecting a collection of links in the direct lattice. A choice of its Poincaré dual is a 1-form which is zero everywhere except on the intersected links, on which it has unit value in the direction normal to the surface. The charge operator, Eq.˜90, then counts (with signs, because $\cramped{\hat{n}_{-c}=-\hat{n}_{c}}$) the number of strings piercing the surface $\tilde{\Sigma}$, i.e. the intersection number between closed strings and the surface. Because the charges are closed strings, this intersection number must be zero unless $\tilde{\Sigma}$ winds around a periodic boundary and intersects a charge which also winds around the periodic boundaries in the transverse direction.
In a closed system, gauge invariance guarantees that the matter charge operators are exactly zero because of the Gauss law, Eq.˜87, which when inserted into Eq.˜90 yields $\cramped{\langle\hat{E},\mathrm{d}\delta_{\tilde{\Sigma}}\rangle=0}$. Equivalently, the intersection numbers, Eq.˜94, are exactly zero because $\partial M$ has trivial homology class, i.e. $\cramped{\delta_{\tilde{\Sigma}}(\partial M)=\mathrm{d}\delta_{\tilde{\Sigma}}(M)=0}$ since $\tilde{\Sigma}$ is closed. Consider for example the case $k=2$, where the charge operators count the number of charged strings wrapped around a periodic boundary. We cannot, in a closed system, have a single non-contractible charged string, because it is attached to an electric membrane which must end somewhere inside the system. The only possibility is that it ends on another non-contractible charged string going the other direction, such that the two strings constitute the boundary of the electric membrane. This is the sense in which a closed system is charge-neutral when $k>1$.
[…]
4.1.2 Open Boundaries and Global Matter Symmetry
To make the charge operators non-trivial, we must impose boundary conditions such that either $\tilde{\Sigma}$ can terminate on the boundary, and $\delta_{\tilde{\Sigma}}$ is not a closed form, or such that the contour $\partial M$ can terminate at the surface and thus not be closed. The latter case, however, means that the charged objects can pass through the boundary, meaning charge is not conserved. Therefore, extending our results from Section˜2, we take the former case. We consider open boundaries in the same form as Fig.˜2. To be precise, the boundary consists of a layer of $D$-dimensional hypercubes, with all cells in the boundary layer which do not touch a bulk cell removed (or, equivalently, $A=0$ and $\theta=0$ on those cells). In other words, we remove the cells on the vacuum side, creating links missing one end, plaquettes missing one edge, cubic cells missing one face, etc. The boundary action is given by